OpenStudy (anonymous):

Armando put $78 into a CD that pays 4.9% interest compounded monthly. According to the Rule of 72, approximately how long will it take for his money to double? A.15.9 years B.14.7 years C.146.9 years D.159.2 years

OpenStudy (anonymous):

? :[

OpenStudy (anonymous):

double of money is 78*2 = 156 then: 156 = 78*(1+0.049/12)^n we need to find n: 156/78 = (1+0.049/12)^n using Log: Log(2) = n*Log(1+0.049/12) Log(2)/Log(1+0.049/12) = n then: n= 170 months

OpenStudy (anonymous):

B 14.7 ?

OpenStudy (anonymous):

14.17 You must have a typing error.

OpenStudy (anonymous):

time = 170 months/12 months per year = 14.17

OpenStudy (anonymous):

14.7 years